Brenda can do a job in 10 days; Adam does the same job in 15 days. After Adam works alone for 3 days, Adam and Brenda work together and finish the job. How many days did Brenda work. (Answer is decimal)
Answer by scott8148(6628)

in the 3 days he works alone , Adam completes 3/15 or 1/5 of the job
___ Adam and Brenda then work together to complete the remaining 4/5
___ Brenda's days are the same as the together days

(t / 10) + (t / 15) = 4/5

multiplying by 30 (LCD) ___ 3t + 2t = 24 ___ 5t = 24